11. Lee owns a houseboat on Lake Tahoe that he personally uses for 12 days out of the year and rents for 280 days. For tax purposes, the houseboat is classified as

A) property that is treated as a hobby which gives rise to from AGI deductions only.

B) a combination of the taxpayer's residence and rental property. The deduction for expenses is limited to the amount of income generated by the property.

C) rental property. Expenses in excess of income may be deducted although net income or loss is subject to the passive activity rules.  

D) neither a residence nor rental property. Because it is rented a nominal number of personal-use days, both revenue and expenses (other than those otherwise allowable) are ignored.

12. Mustafa owns a trucking business. During the current year he incurred the following:

Gasoline and oil

$ 75,000

Maintenance

$ 18,000

Fines for speeding

$ 2,500

Bribes to government inspection officials

$ 7,500

The fines for speeding were a necessary cost because missing deadlines would cause lost business and are ordinary in the industry. What is the total amount of deductible expenses?

A) $28,000

B) $75,000  

C) $85,000

D) $93,000

13. Katarina owns a condominium in the Hilton Head Island, South Carolina. During the year, Katarina uses the condo a total of 30 days. The condo is also rented to vacationers for a total of 150 days and generates rental income of $20,000. Katarina incurs the following expenses:

Expense

Amount

Mortgage interest

$ 7,500

Property taxes

4,800

Utilities

1,800

Insurance

1,200

Depreciation

9,000

Using the IRS method of allocating expenses, the amount of depreciation that Katarina may take with respect to the rental property will be

A) $ 1,050.

B) $ 6,780

C) $ 7,250.

D) $9,000.
Answer #1

Answer 11.

Statement c is correct (as being used for mostly rental purposes and not for own use)

Answer 12.

Deductible expenses = 75000 + 18000 = 93000

Correct option is D

Answer 13.

Net rental income = 20000 - [(7500+4800+1800+1200) * 150/180 = 20000 - 12750 = 7250

Depreciation to be charged = 9000 * 150/180 = 7500

Depreciation is more than rental income, thus

Depreciation can be deducted only upto $7250

Correct option is C